# IRIF, Université Paris Cité, France

### v1

@@ 4.1 Identifying Deeper Merges
While going through the signature extraction procedure, we made the crucial assumption that $x$ is a critical point of a neuron on the layer we are targeting. Since we are trying to extract that layer, we cannot verify this assumption. We might be extracting a neuron of a deeper layer. The authors of [8] claim that it is exceedingly unlikely that signatures extracted from critical points on a deeper layer can be merged into a component. They conclude that a component of size greater than two is on the target layer. In this section, we explain why signatures from deeper layers can indeed be merged, discuss why this can be a problem, and propose potential solutions to address this issue.
@@
When we extract from two critical points $x_{1}$ and $x_{2}$ of the same neuron in a deeper layer $i+t$ using its preceding layer $i+t-1$, we solve for $y_{1}$ and $y_{2}$ in the respective systems below and observe that $y_{1}$ and $y_{2}$ can be merged:
@@
Let’s write $A_{x_{1}}$ the matrix $\ell^{(i+t-1)}\circ I^{(i+t-2)}_{x_{1}}\circ\dots\circ I^{(i)}_{x_{1}}\circ\ell^{(i)}$ and $A_{x_{2}}$ the matrix $\ell^{(i+t-1)}\circ I^{(i+t-2)}_{x_{2}}\circ\dots\circ I^{(i)}_{x_{2}}\circ\ell^{(i)}$. Thus, we can rewrite these two systems with $A_{x_{1}}$ and $A_{x_{2}}$,
@@
Now suppose that $x_{1}$ and $x_{2}$ have the same activation pattern - meaning they set the same neurons as active neurons per layer - between layers $i$ and $i+t-1$. In this case, $A_{x_{2}}=A_{x_{1}}$, which we write as $A$. Therefore, our systems are as follows.
@@
How does this noise behave? First, the smaller the width of the network, the larger the noise, as critical points from deeper neurons are partitioned into fewer components. Second, the deeper the network, the larger the noise, as critical points on the target layer represent a smaller fraction of the critical points we find. We believe this is one of the reasons why previous works still managed to perform the extraction on shallow networks.
In addition, in both [8] and [13]’s implementations, the selected neurons are the $d_{i}$ largest components of size greater than two. When using this criterion, noise becomes a problem when the size of a noise component becomes larger than the size of a component corresponding to a neuron on the target layer.
We illustrate the issue of deeper critical points with an example slightly deeper than what previous works managed to recover: using $3,000$ critical points we extract the 4th layer of a model trained on the MNIST dataset with architecture $1784-8^{(8)}-$ 1. Due to space constraints, only the largest 16 components are displayed. As shown in Figure 4, four of the first eight components recovered by the original signature extraction belong to deeper layers. This shows that, in deeper networks, the size of a component is not a sufficient indicator to identify if the neuron is on the target layer.
Figure 4: Merging results from the original attack on layer 4 of a $1784-8^{(8)}-$ 1 MNIST model with $3,000$ critical points. Each component on the target layer is labelled with its associated neuron on top. Larger component size does not necessarily indicate that the component belongs to the target network.
Discarding points on deeper layers. To discard points on deeper layers, we first recycle and improve an algorithm from [8] that was only used in the context of the targeted search. The process involves 1) computing 100 distinct intersection points between the extracted hyperplane and the extracted network for the critical point under test, and 2) verifying on the target neural network whether these points lie on the extracted hyperplane. One intersection point not belonging to the extracted hyperplane indicates that the hyperplane has bent on a neuron that has not yet been extracted, and hence that the extracted hyperplane is on a deeper layer. This test is not an if-and-only-if condition, as a hyperplane not breaking does not guarantee that it corresponds to a neuron on the target layer (see Figure 6). We apply this test to every critical point we find. This reduces the noise, but it is not sufficient. As exemplified in Figure 5, though we remove one component from deeper layers, four large such components remain.
Figure 5: Merging results after discarding critical points on deeper layers. Same experimental setup as in Figure 4. Components on deeper layers either disappear or have a smaller size.
Discarding deeper components. We need to introduce additional discrimination techniques to reduce the interference of deeper layers. While the strategy above aimed at discarding points on deeper layers, we now turn to components. We use three criteria. First, we know that critical points that merge usually belong to the same neuron, regardless of whether they are in the target layer or a deeper layer. Therefore, the more critical points identified from deeper layers by the above test that merge with our candidate component, the more likely it is that this component belongs to a deeper layer. For this reason, we compute for each component a noise ratio
@@
Second, because merges of deeper critical points come from critical points with the same partial activation pattern, those deeper components have a harder time obtaining a diverse set of critical points to yield a neuron with a very small number of entries unrecovered. Finally, because of the constraint on the activation pattern of deeper merges, components on deeper layers tend to be smaller in size. While none of these criteria are if-and-only-if, combining them yields satisfying results. First, we remove components with a size smaller than a certain fraction of the largest component (we found that in practice, $0.1$ was sufficient for networks with eight hidden layers). Second, we discard components that either have a $\tau$ above a certain threshold (we use $0.1$ for networks with width 8 and $0.2$ for networks with width 16) or a large number of entries unrecovered (at least half of them) if their $\tau>0$.
It’s worth noting that making a strict association between deeper critical points and incorrect components could lead to mislabelling. For example, critical points of a neuron in a deeper layer $i+t$ can merge with those of neuron $\eta_{k}^{(i)}$ on the target layer $i$, if they cause $\eta_{k}^{(i)}$ to be the only active neuron on the target layer (proof in Appendix 0.D).
On a side note, when this happens, the sign of all neurons on layer $i$ can be recovered immediately with no queries or computations as we have an input $x^{*}$ for which all neurons on the layer must be inactive (first order derivatives in $\mathcal{P}_{x^{*}}$ are equal to zero). We must have $\eta^{(i)}_{k}\cdot x^{*}+b^{(i)}_{j}<0$, for all neurons on the layer, so we choose the sign of each $\eta^{(i)}_{k}$ to match this condition.
The graph illustrating the effectiveness of discarding deeper components is presented later in Figure 9, as the attack first needs to deal with another kind of limitation.
$\eta_{1}^{(i-1)}$$\eta_{2}^{(i-1)}$$\eta^{(i+j)}$$\eta^{(i+k)}$$x$$a_{2}$$a_{1}$
Figure 6: Identifying if $x$ is on the target layer. $0\leq j<k$. By finding that $a_{2}$ is not a critical point, we can infer that the hyperplane we extracted from $x$ (- - -) broke on a layer $i+j$ we have not extracted (—-). Therefore, $x$ cannot be on layer $i$. However, finding that $a_{1}$ is a critical point does not give any information regarding $x$’s layer.
4.2 Increasing Rank with Subspace Intersections
When extracting the signature, our goal is to obtain a partial signature $y_{x}$ up to an unknown constant from a critical point $x$ (see Section 3.4). As a result, the solution space should be of dimension 1. To determine $y_{x}$, we solved the equation $(\Gamma_{x}^{(i-1)}\Delta)\cdot y_{x}=\partial^{2}_{\Delta}f(x)$ by selecting a sufficient number of random directions $\Delta$. The partial row obtained consists of coefficients corresponding to the active neurons in layer $i-1$ for the given input $x$. We denote by:
  • $s_{x}^{(i-1)}$: the number of active neurons in layer $i-1$,
  • $r_{x}^{(i-1)}:=\operatorname{rank}(\Gamma_{x}^{(i-1)})$.
If $s_{x}^{(i-1)}>r_{x}^{(i-1)}$, the system is underdetermined and cannot be solved. This issue arises when there are fewer active neurons in a previous layer $k<i-1$ than in layer $i-1$ (see Figure 7), as indeed:
$\Gamma_{x}^{(i-1)}=I_{x}^{(i-1)}A^{(i-1)}\cdots I_{x}^{(1)}A^{(1)}$
and thus,
$\text{rank}(\Gamma_{x}^{(i-1)})=\min(\{\text{rank}(I_{x}^{(k)})\}_{1\leq k\leq i-1})\neq\text{rank}(I_{x}^{(i-1)})$
If $X$ denotes the random variable representing the number of active neurons per layer, then $\mathbb{P}(r_{x}^{(i-1)}<s_{x}^{(i-1)})=1-\mathbb{P}(X\geq s_{x}^{(i-1)})^{i-1}$. Because of the $i-1$ exponent, it appears infeasible to extract deeper layers beyond the first few.
@@ 4.2 Increasing Rank with Subspace Intersections
Figure 7: Solving for a neuron’s weights with certain critical points can lead to underdetermined systems. Target neuron is in black. Active neurons are in blue. The rank of $\Gamma_{x}^{(1)}$ is 2, which limits the rank of $\Gamma_{x}^{(2)}$ to 2 while we expect 3.
Signature intersections. If the rank is insufficient, we no longer have a unique solution but rather a set of solutions. For a critical point $x$ in the layer $i$, the solution space is given by $\mathcal{S}_{x}=L_{x}+\ker(\Gamma_{x}^{(i-1)})$, where $L_{x}$ is a solution of the solution space and $\ker(M)$ represents the kernel space of a matrix $M$. In [8], a method was proposed to merge two partial signatures. Here, we extend this approach to merge two solution spaces.
Suppose we have another critical point $x^{\prime}$ with solution space $\mathcal{S}_{x^{\prime}}=L_{x^{\prime}}+\ker(\Gamma_{x^{\prime}}^{(i-1)}).$ Since we recover rows up to a scalar factor, we seek to compute the intersection of $\mathcal{S}_{x}$ and $\lambda\mathcal{S}_{x^{\prime}}$, where $\lambda$ is a scalar factor equal to the ratio of the row scalar factors of $x$ and $x^{\prime}$. Let $(e_{1},\ldots,e_{k})$ be a basis of $\ker(\Gamma_{x}^{(i-1)})$, and let $(f_{1},\ldots,f_{r})$ be a basis of $\ker(\Gamma_{x^{\prime}}^{(i-1)})$. We then solve the following system in $\mathbb{R}^{d_{i-1}}$:
$L_{x}+\mu_{1}e_{1}+\dots+\mu_{k}e_{k}=\lambda L_{x^{\prime}}+\lambda_{1}f_{1}+\dots+\lambda_{r}f_{r},$
where the unknowns are $\mu_{1},\dots,\mu_{k},\lambda,\lambda_{1},\dots,\lambda_{r}\in\mathbb{R}$. This system consists of $d_{i-1}$ equations in $\mathbb{R}$. However, only the equations corresponding to the active neurons in the layer $i-1$ for both inputs $x$ and $x^{\prime}$ are relevant. To ensure that we do not merge spaces from critical points associated with different rows, we overdetermine the system. Indeed, an overdetermined system with random coefficients is inconsistent with very high probability, and thus a system of $\ell>N$ equations with $N$ unknowns typically doesn’t have a solution except if there is some ground truth behind it. Therefore, we impose the following merging condition:
$\ell>1+|\ker(\Gamma_{x}^{(i-1)})|+|\ker(\Gamma_{x^{\prime}}^{(i-1)})|=1+k+r,$
where $\ell$ is the number of relevant equations.
Intersections allows us to exploit at least 90% of critical points on the target layer when attacking the $1784-8^{(8)}-$ 1 MNIST model. As a comparison, the original attack would have solved a system with adequate rank and extracted a correct signature on only 6% of the critical points of layer 4. We could intersect subspaces three by three to have a lower condition for merging, allowing for the use of more critical points. However, merging three by three instead of two by two slows down considerably the attack. For this reason, all our results use two-by-two merges.
As shown in Figure 8, after intersecting critical points with insufficient ranks, the top $7$ components yield correct signatures that correspond to neurons in the target layer. Then, after applying our additional discrimination techniques described in Section 4.1 to discard deeper components, all remaining components are in the target layer, as shown in Figure 9. The only unrecovered component corresponds to neuron $\eta^{(4)}_{5}$, which is always-off (see Figure 10).
Figure 8: Merging results after signature intersections. Components with rank issues are discarded, and correct components become prominent. Neuron $\eta_{5}^{(4)}$ is always off (see Figure 10).
Figure 9: Final merging results after discarding deeper components. All remaining components are components on the target layer without rank issues. Neuron $\eta_{5}^{(4)}$ is always off (see Figure 10).
5 Experiments
All our experiments were conducted on an Intel Core i7-14700F CPU using networks trained on either the MNIST or CIFAR-10 datasets, two widely used benchmarking datasets in computer vision. They have also been commonly used in related works [8, 6, 13] to evaluate model extraction methods. MNIST consists of 28 $\times$ 28 pixel grayscale images of handwritten digits, divided into ten classes (“0” through “9”) [17]. In contrast, CIFAR-10 contains 32 $\times$ 32 pixel RGB images of real-world objects across ten classes (e.g., airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck) [16].
We evaluate our attack on three different networks: one trained on CIFAR-10 with architecture $-13072\text{-}8^{(8)}\text{-}$ 1 and two networks trained on MNIST with architectures $-1784\text{-}8^{(8)}\text{-}$ 1 and 784-$-1784\text{-}16^{(8)}\text{-}$ 1. The main results are given in Table LABEL:table:result_summary_table. We refer to those models as Model I, II, and III, respectively. Following [13], we extract the network layer by layer, each time assuming that the previous layers have been correctly recovered. This crucial assumption allows us to evaluate the effectiveness of our improvements, setting aside the floating-point computation imprecisions. State-of-the-art extraction from [13] could not extract any hidden layer past the third layer on a $1784-16^{(8)}-$ 1 MNIST trained network. When attacking the same architecture, we can recover over 95% of the weights in each deeper layer. We apply the attack to other network architectures, yielding similar results. Our improvements largely outperform existing signature extraction methods (Table LABEL:table:result_summary_table).
We attempted to implement end-to-end model extraction without assuming perfect recovery of the preceding layers, instead relying on previously recovered results to extract subsequent layers. However, for Model II, accumulated floating-point imprecision errors emerged by Layer 5, resulting in incorrect partial signatures. For the same reason, the extraction of Models I and III stopped at layer 3. We also tried to attack a much deeper network with architecture $1784-8^{(16)}-$ 1. The algorithm to discard critical points becomes less effective. This makes deeper components difficult to distinguish based on either $\tau$ or size, and intersection computation times grow significantly. Furthermore, because our networks only have width 8 for runtime purposes, when we reach the final four layers (layers 13–16), the solution spaces of critical points almost always have rank $1$. The merging condition is rarely met, and we recover no components. We believe this should not happen when attacking wider networks. Therefore we focused on networks with eight hidden layers as those are the deepest considered in the literature of both signature [13] and sign recovery [6].

### v2

@@ 3.2 Deeper Layers’ Influence on Extraction
While going through the signature extraction procedure, we made the crucial assumption that $x$ is a critical point of a neuron on the layer we are targeting. Since we are trying to extract that layer, we cannot verify this assumption. We might be extracting a neuron on a deeper layer. The authors of [8] claim that it is exceedingly unlikely that signatures extracted from critical points on a deeper layer can be merged into a component. They conclude that a component of size greater than two is on the target layer. In this section, we first explain why signatures from deeper layers can indeed be merged. Second, we propose a two-step solution which consists in discarding deeper points, then deeper components. Finally we discuss how noise behaves in very large networks.
@@
The issue: deeper merges. Assume that we are trying to extract layer $i$. From two critical points $x_{1}$ and $x_{2}$ of the same neuron in a deeper layer $i+t$, we solve respectively for partial weights $y_{1}$ and $y_{2}$ using the preceding layer i+t1i+t-1 in the systems below. Since $x_{1}$ and $x_{2}$ are critical points of the same neuron, and we are extracting from the preceding layer, we know that $y_{1}$ and $y_{2}$ can be merged. We’ll see that under a specific scenario, the extraction from $x_{1}$ and $x_{2}$ can also be merged when using layer i1i-1 for extraction, inducing noise in our extraction.
@@
Let’s write $A_{x_{1}}$ the matrix $I^{(i+t-1)}_{x_{1}}\circ A^{(i+t-1)}\circ I^{(i+t-2)}_{x_{1}}\circ\dots\circ I^{(i)}_{x_{1}}\circ A^{(i)}$ and $A_{x_{2}}$ the matrix $I^{(i+t-1)}_{x_{2}}\circ A^{(i+t-1)}\circ I^{(i+t-2)}_{x_{2}}\circ\dots\circ I^{(i)}_{x_{2}}\circ A^{(i)}$. Thus, using $A_{x_{1}}$ and $A_{x_{2}}$, we can rewrite
@@
Now suppose that $x_{1}$ and $x_{2}$ have the same activation pattern between layers $i$ and i+t1i+t-1, meaning they set the same neurons as active. In this case, $A_{x_{2}}=A_{x_{1}}$, which we write as $A$. Therefore, our systems are as follows.
@@
[…]
Figure 5: Identifying if $x$ is on the target layer. $0\leq j<k$. By finding that the intersection point $a_{2}$ is not on the extracted hyperplane, we infer that the hyperplane we extracted from $x$ (- - -) broke on a layer $i+j$ we did not extract (). Thus, $x$ cannot be on layer $i$. Yet, finding that $a_{1}$ is on the extracted hyperplane does not give any information about $x$’s layer.
Discarding deeper points. To discard points on deeper layers, we recycle and improve an algorithm from Section 4.4.2 of [8] that was only used in the context of the targeted search. For each critical point found, the algorithm process involves: 1) computing a large number of distinct intersection points (we use 100 in practice) between the hyperplane extracted from the critical point under evaluation and the extracted network, and 2) verifying on the target neural network whether these intersection points still lie on the extracted hyperplane. One intersection point not belonging to the extracted hyperplane indicates that the hyperplane has bent on a neuron that has not yet been extracted, and hence that the extracted hyperplane is on a deeper layer. This test is not an if-and-only-if condition, as a hyperplane not breaking does not guarantee that it corresponds to a neuron on the target layer, see Fig. 5. This method reduces the noise, but it is far from sufficient, see Fig. 3(b). We refer the reader to the original description for a more thorough account of the strategy.
Discarding deeper components.
After discarding points on deeper layers, we compute intersections as described in Section 3.1, see Fig. 2. Intersections deal with components having insufficient rank. However, some components on deeper layers remain, see Fig. 3(c). We use three criteria to discard them. First, we know that critical points that merge usually belong to the same neuron, regardless of whether they are in the target layer or a deeper layer. Therefore, the more critical points identified from deeper layers by the above test that merge with our candidate component, the more likely it is that this component belongs to a deeper layer. For this reason, we compute for each component a noise ratio:
@@
Second, because merges of deeper critical points come from critical points with the same partial activation pattern, deeper components have a harder time obtaining a diverse set of critical points to yield a neuron with a very small number of entries unrecovered. Finally, because of the constraint on the activation pattern of deeper merges, deeper components tend to be smaller in size. While none of these criteria is an if-and-only-if test, combining them yields satisfying results.
Therefore, we first remove components with a size smaller than a fraction of the largest component (we found that in practice, $0.1$ was sufficient for networks with eight hidden layers). Then, we discard components that either have a $\tau$ above a certain threshold (we use $0.1$ and $0.2$ for networks with width $8$ and $16$, respectively) or a large number of entries unrecovered (at least half of them) if $\tau>0$. Notably, making a strict association between deeper critical points and incorrect components can lead to mislabelling (example in Appendix 0.D).
Noise in large networks. We have shown above that components coming from deeper layers result from critical points with the same activation pattern between the target layer and their layer. As the width of the network increases, the neurons on each layer cut the input space into smaller polytopes, making it more unlikely that critical points share the same partial activation pattern, up to the point of noise not being an issue anymore for a large enough width. For example, when extracting layer $4$ of a $784-256^{(16)}-1$ network with $100,000$ critical points, we get 189 components of size larger than $2$. Only three of these do not belong to the target layer, with sizes $3$, 33, and $4$. By contrast, $177$ components have size at least $5$, and $157$ have size at least $20$. This shows that, in wide networks, imposing a minimum component size is an effective way to filter out noise.
As network width grows, intersection computations become a problem. With width $256$, each critical point gives many recovered weights. This allows previously unavailable statistical arguments by viewing recovered weights as realisations of random variables. Indeed, the contribution of a neuron on a deeper layer corresponds to a very different mathematical expression from a neuron on the target layer. The deeper the neuron we are actually extracting upon, the higher the expected variance of the recovered weights is. Therefore, recovered neurons with unusual high variance can immediately be deemed to be on a deeper layer, without requiring any queries or computation, thereby increasing the proportion of critical points on the target layer. The time and queries for finding critical points grows linearly with the number of points we require, while computing intersections is comparatively very slow and grows polynomially with the number of points processed. For this reason increasing the proportion of critical points on the target layer, even if we discard some of them, results in a considerable speed-up of the attack. See Appendix 0.E for proofs, experiments and more details.
3.3 Numerical Precision in the Context of an End-to-End Attack
In practical end-to-end attacks, model extraction relies on finite-precision arithmetic (typically 64-bit floating-point). However, many algorithms in the extraction process are numerically unstable, introducing small errors in the recovered weights. Such errors accumulate across layers and block the extraction of deeper layers. Thus, it is crucial to refine the recovered weights to minimise imprecision.
Carlini et al. [8] introduce a method to improve the precision of recovered signatures. It consists in finding more critical points, at least $d_{i-1}$ (the number of neurons in layer i1i-1) for each neuron $\eta_{k}^{(i)}$ in the target layer $i$, and then solving the following linear system for its weight vector $\hat{A}^{(i)}_{k}$ and bias $\hat{b}^{(i)}_{k}$:
@@ 3.3 Numerical Precision in the Context of an End-to-End Attack
where $h_{j}=\hat{F}^{(i-1)}(x_{j})$ and $x_{j}$ is a critical point for neuron $\eta_{k}^{(i)}$. This leads to precision improvements as weights given by components come from multiple imprecision-inducing computations (derivatives, system solving, etc.) which themselves depend on the precise identification of critical points.
To find more critical points, Carlini et al. [8] use a kind of targeted search. Foerster et al. [12] note that this search often struggles in practice and propose refinements, but these add numerous parameters hindering the reproducibility and substantially increasing the runtime. They report precision-improvement runtimes of 33 times that of signature extraction and 17 times that of sign extraction. For this reason, [12] recommends avoiding precision improvements, while noting that it is a “remaining concern” for an end-to-end attack.
Our improved signature extraction can efficiently avoid this issue. Specifically, we explicitly generate a larger pool of critical points at the beginning, a number far exceeding that required in Carlini et al. [8]. These points primarily serve to filter deep points and components in the early extraction steps, and are sufficient to be reused for precision refinement, eliminating the need for a dedicated, parameter-heavy targeted search. Our precision improvement step therefore remains a small fraction of the total extraction.
Carlini et al.’s precision improvement [8] is very potent but can only be applied at the end of the extraction of each layer. In other words, the attack per layer is still ran on imprecise numbers. Let’s see why this can be a problem, before presenting our improvement to Carlini et al.’s method for wide layers.
Imprecision due to scaling factors affects extraction efficiency. When recovering a partial weight-vector $y_{x}$ at a critical point $x$ from a linear equation system $(\hat{\Gamma}_{x}^{(i-1)}\Delta)\cdot y_{x}=\partial^{2}_{\Delta}f(x)$, each output from the previous layer is scaled by a neuron-specific factor, denoted as $(c_{1}^{(i)},c_{2}^{(i)},\cdots,c_{d_{i}}^{(i)})$. Large disparities among these scaling factors can make the matrix $\hat{\Gamma}_{x}^{(i-1)}\Delta$ ill-conditioned, amplifying numerical errors so that even tiny errors in recovered weights $\hat{\Gamma}_{x}^{(i-1)}$ cause large deviations in the partial weight-vector $y_{x}$. Critical points with such imprecise $y_{x}$ cannot be merged into components, forcing the collection of more critical points initially to assemble a component capable of recovering all desired weights. As discussed in Section 3.2, computing intersections is already slow and grows polynomially with the number of points processed; imprecise partial weights further increase the number of queries and the cost of signature intersections, reducing overall efficiency.
To address this, we normalize each column of $\hat{\Gamma}_{x}^{(i-1)}\Delta$ by multiplying it by a secondary scaling factor $(c_{1}^{(i)^{\prime}},c_{2}^{(i)^{\prime}},\cdots,c_{d_{i}}^{(i)^{\prime}})$, bringing all elements within [1,1][-1,1]. After solving the normalized linear equation system, these factors are multiplied back onto the solution to obtain the recovered signature. This normalization ensures more stable and precise signature extraction, see Appendix 0.F for an example. For instance, when our normalization method is applied to layer 3 of Model II, the precision of partial weights improves by roughly 10 times. The proportion of critical points merged into components increases from $88.9\%$ to $96.2\%$, capturing nearly all points associated with the target-layer neurons.
Imprecision from critical points in large networks. We identify critical points by computing expected intersections of lines given by output derivatives, see Appendix 0.B. The search for critical points can therefore be imprecise if the change in linearity caused by a critical hyperplane is small, i.e. if the output region is rather flat, see Fig. 6 for illustration. We thus classify critical points based on the angles between the second derivatives used to identify them. We then only stack up critical points with large derivative differences to construct the matrix in ($\star$3.3). For networks with a small width of 8, the number of flat critical points originally used in ($\star$3.3) is small and does not affect much the precision of the recovered neurons. Yet, we notice significant improvement for larger widths, for example a $\times 3$ precision improvement on the crucial first layer of our Model II.
$xx^{*}$
$^{*}xx$
Figure 6: Illustration of imprecisions in flat regions when finding critical points. In both cases the attacker computes a slightly imprecise derivative (dashed red line). Though the angular error (angle between the dashed red line and the black line next to it) is the same in the two cases, the distance between the extracted critical point $x^{*}$ and the real one $x$ changes depending on how flat the region is. The worst errors are caused by flat regions. For this reason we avoid them.
4 Confident Sign Extraction
In their original presentation of the neuron wiggle, the authors of [6] managed to recover almost all neuron signs for a $3072-256^{(8)}-10$ neural network, assuming all signatures had been correctly recovered. It failed for neurons with so-called low confidence. This is why the authors of [6] recommend using more resources to ensure the sign extraction of the $10\%$ lowest-confidence neurons. However, the authors of [12] found that in practice, on smaller architectures, many neurons have low confidence and, worse, increasing resources does not improve it. They resort to an exhaustive search over the signs of the low-confidence neurons to ensure successful recovery. Indeed, making a mistake on the sign of a neuron typically leads to a failure of the signature recovery of the following layer. This calls for a polynomial method to find the sign of low-confidence neurons.
We propose to combine SOE and the neuron wiggle to obtain a sign-extraction strategy which uses neurons with higher confidence than the one used by the neuron wiggle alone. We set-up the SOE under-determined system at a point $x$, and then use the result of the neuron wiggle on high-confidence neurons to complete the rank: we choose the neuron inactive at $x$ with the highest confidence, we eliminate its variable from the SOE unknowns. We iteratively add the most confident inactive neurons until the system reaches sufficient rank. Solving this system then gives the signs of the remaining low-confidence inactive neurons at $x$ alongside all the active neurons, overcoming the limitations of the neuron wiggle. For this strategy to succeed, we need a point $x$ where the number of neurons on the target layer inactive at $x$ is at least $d_{i}$ minus the SOE rank. We found such points easily for all networks considered by sending random input points through the extracted network. We also give a way to increase the rank of SOE. This rank-increasing method pushes further the value of combining SOE and the neuron wiggle, increasing the confidence of sign recovery.
As explained in Section 2.4, SOE consists in solving at a point $x$ and for multiple directions $\Delta_{k}$ the system,
$\{G_{x}^{(i+1)}I_{x}A^{(i)}F^{(i-1)}\Delta_{k}=f(x+\Delta_{k})-f(x)\}_{k},$
where the unknown is $G_{x}^{(i+1)}I_{x}$. The rank of the system is limited by the rank of $F^{(i-1)}$ around $x$. Crucially, because the unknown is $G_{x}^{(i+1)}I_{x}$, any critical point with the same later activations (i.e., the same $G^{(i+1)}I^{(i)}$) can be used even if the previous activation patterns differ. This implies that we do not have to build the system around a single critical point. Instead, we search for different linear regions nearby which keep later activations fixed. This triggers many different activation patterns for $F^{(i-1)}$, their union having a higher rank, see Fig. 7. More details are given in Appendix 0.G. We call this strategy of building the SOE system not from a single critical point but from a large zone an extension of SOE.
$\eta^{(i+j)}$$\eta_{1}^{(1)}$$\eta_{2}^{(1)}$$\eta_{1}^{(2)}$$\eta_{2}^{(2)}$
Figure 7: Identifying critical points that share the same activation pattern after the target layer $i$. First, we randomly select a line () in the input space and identify the largest continuous segment of critical points (•) from previous layers that lies between two critical points () in deeper layers. Although these points differ in activation patterns in the previous layers, they share the same activation pattern in the deeper layers. We then take the midpoint () of this segment as the center of a sphere and explore multiple random directions (- - -) to find more critical points from previous layers within this region.
This yields two different strategies depending whether we favour a large zone to extend to, or an initial point with a high SOE rank and a large number of confident inactive neurons. Both strategies start by computing the neuron wiggle for all neurons. If we favour the former, the extension strategy, we first run random lines through the input space, then we perform the extension on the most promising zones (see Fig. 7). We then complete with the highest confidence neurons and select the zone with the highest lowest confidence used. The earlier we are in the extraction, the smaller the equivalence classes defined by $x_{1}\sim x_{2}\iff\{G^{(i+1)}_{x_{1}}I_{x_{1}}=G^{(i+1)}_{x_{2}}I_{x_{2}}\}$ are, and the less effective this increase in the rank of SOE is. This strategy works at its best in the deeper layers. If we favour the latter, the confidence strategy, we run random points through the network, compute the minimum number of active neurons per layer (the rank of SOE) and we select the point which maximises the lowest confidence used to solve the signs. We then perform an extension on that point before solving for the sign. This strategy works at its best in the first layers where the rank of SOE is still high. The authors of [6] argue that the rank of SOE stabilises in the deeper layers. Surprisingly, both strategies gave similar results in our experiments though they rely on two different structural properties of the target network. The extension strategy relies on the number of previous layer critical points within possible extensions whereas the confidence strategy relies on the rank of SOE as we progress through the network. An attacker can run both strategies and pick for each layer the one giving the more confident results. Variants between the two strategies can also be done (extension on more points than the $x$ given by the confident strategy, assessing the SOE rank of the midpoint of the extension strategy etc.). We refer to all these new sign recovery strategies as SOE+Wiggle. For the sake of consistency, unless stated, all our experiments use the extension strategy. This is because for networks our end-to-end attack targets, the extension strategy can completely avoid the neuron wiggle for many layers, giving a deterministic method for sign recovery.
We compare our method with prior work on Model II in Table 1. Using the neuron wiggle alone, signs can be recovered only in layers 6 and 7; in these layers, the confidence falls below the 0.75 minimum threshold recommended by [12]. In contrast, SOE+Wiggle recovers all signs correctly across the network with high confidence and without resorting to exhaustive search. For large-scale networks, SOE+Wiggle remains effective. In Table 2, we compare our method with prior work on the $3072-256^{(8)}-10$ model used in Table 4 of [6]. When computing each neuron’s confidence using the neuron wiggle, we limit the total budget to $10^{6}$ critical points, instead of collecting 200 points per neuron as in [6]. We set this limited budget because some neurons would require an excessive number of queries to obtain 200 critical points. With this budget, we still ensure that every neuron has a number of critical points that exceeds the threshold recommended in [12] for stable confidence estimates. Using the neuron wiggle alone, some signs in layers 5, 7, and 8 are incorrectly recovered, and the lowest confidence in each layer is low (around its minimum value 0.5). In contrast, our SOE+Wiggle method recovers all signs correctly across the network with high confidence.
Table 1: Comparison of sign-extraction methods on Model II. When recovering the signs of neurons on the target layer, we assume that all preceding weights have been correctly extracted.
Method Metric L2 L3 L4 L5 L6 L7 L8
Neuron Wiggle [6] correct sign extraction $7/8$ $5/8$ $5/8$ $6/8$ $8/8$ $8/8$ $5/8$
Neuron Wiggle + exhaustive search [12] low-confidence neurons ($\alpha<0.75$) $4/8$ $5/8$ $5/8$ $4/8$ $4/8$ $3/8$ $4/8$
SOE + Wiggle correct sign extraction $8/8$ $8/8$ $8/8$ $8/8$ $8/8$ $8/8$ $8/8$
(our method) lowest confidence used $\infty^{\dagger}$ $0.81$ $\infty$ $\infty$ $0.75$ $1.0$ $\infty$
L stands for layer. When the SOE matrix reaches full rank after extension, the sign extraction is deterministic and the neuron wiggle is unnecessary.
Table 2: Comparison of different sign-extraction methods on the $3072-256^{(8)}-10$ model used in Table 4 of [6]. When recovering the signs of neurons on the target layer, we assume that all preceding weights have been correctly extracted.
Method Metric CIFAR10-256(8)-10
L2 L3 L4 L5 L6 L7 L8 Recovery failures
Neuron Wiggle [6] lowest confidence used 0.53 0.53 0.51 0.50 0.50 0.51 0.51 15
Confidence strategy lowest confidence used 0.87 0.83 0.75 0.67 0.64 0.62 0.78 0
confidence increases from ext. 0.05 0.08 0.07 0.19
Extension strategy lowest confidence used 0.86 0.80 0.64 0.62 0.56 0.66 0.76 0
confidence increases from ext. 0.09 $\infty^{\ddagger}$ $\infty$ 0.15 $\infty$
L stands for layer. ext. stands for the extension of SOE. When the SOE extension does not increase the SOE rank, the lowest confidence level used remains unchanged. Without SOE extension, the extension strategy fails as the SOE rank plus the number of inactive number at the point $x$ is still smaller than $d_{i}$.
5 Evaluation Metrics and Unrecovered Weights